How can I solve this by elimination method?
x-y=8
6+3y=-2x
And can I solve this in substitution method and get the answer x=18/5 and y=10/5

Answer by stanbon(75887)

How can I solve this by elimination method?
1st: x-y=8
2nd: 6+3y=-2x
------------
Multiply thru 1st by 3 to get:
3rd: 3x-3y=24
2nd: 2x+3y=-6
------------
Add 2nd and 3rd to get:
5x = 18
x = 18/5
-------------
Substitute that into x-y=8 to solve for y:
18/5 -y = 8
y = 18/5 -40/5
y = -22/5
======================

And can I solve this in substitution method and get the answer x=18/5 and y=10/5
No, not if the work above is correct.
==================
Cheers,
Stan H.